What is the equation of 2x+3=4
Found 2 solutions by rothauserc, Mangcha:
Answer by rothauserc(4718)

2x+3=4
2x+3-3=4-3
2x = 1
x = 1/2

Answer by Mangcha(2)

Soln: 2x+3=4
=>2x+3-3=4-3 [subtracting both sides by 3]
=>2x=1
=>x=1/2